Two balls are drawn in succession without replacement from an urn containing 6 green balls and 6 yellow balls. Let Y be the random variable representing the number of yellow balls. Find the values of the random variable Y


1
Expert's answer
2022-01-16T08:54:29-0500

"P(Y=0)=\frac{6}{12}*\frac{5}{11}=\frac{5}{22}."

"P(Y=1)=\frac{6}{12}*\frac{6}{11}+\frac{6}{12}*\frac{6}{11}=\frac{6}{11}."

"P(Y=2)=\frac{6}{12}*\frac{5}{11}=\frac{5}{22}."